OpenStudy (anonymous):

Little jimmy is building a corner cubourd to go in his bedroom and he wants it to run 10cm along each wall. a) use pythagoras theorem to find the exact length of timber required to complete the triangle. b) find a rational approximation for the length, rounding your answer to the nearest millimetre

OpenStudy (anonymous):

Is this question hard?

OpenStudy (anonymous):

Pythagoras theorem is a^2+b^2=c^2 and we know a=b so then solve for c and add each side to find the perimeter. C=200^.5 now add that to 20 and you get 34.14 that's a. b. a rational approximation is what you get from the equation (a-(m/n))<or=1/n i'm not sure how to work that equation though.

OpenStudy (anonymous):

for b i would guess 14.1 because that is the length of c but also it is rounded to the nearest mm

OpenStudy (anonymous):

and no it's not that hard.

OpenStudy (anonymous):

ok ty